# NRGBoost: Energy-Based Generative Boosted Trees

### v1

@@ 3.3 Regularization
The simplest way to regularize a boosting model is to stop training when overfitting is detected by monitoring a suitable performance metric on a validation set. For NRGBoost this could be the increase in log-likelihood at each boosting round. However, estimating this quantity would require drawing additional validation samples from the model (see Appendix A.1). An alternative viable validation strategy which needs no additional samples is to simply monitor a discriminative performance metric. This amounts to monitoring the quality of $q_{f}(x_{i}|\mathbf{x}_{-i})$ instead of the full $q_{f}(\mathbf{x})$.
Besides early stopping, the decision trees themselves can be regularized by limiting the depth or total number of leaves in each tree. Additionally, we can rely on other strategies such as disregarding splits that would result in a leaf with too little training data, $P(X)$, model data, $Q_{f}(X)$, volume $V(X)$ or too high of a ratio between training and model data $\nicefrac{{P(X)}}{{Q_{f}(X)}}$. We found the latter to be the most effective of these, not only yielding better generalization performance than other approaches, but also having the added benefit of allowing us to lower bound the acceptance probability of our rejection sampling scheme. Furthermore, as we show in Appendix A.2, limiting this ratio guarantees that a small enough step size produces an increase in likelihood at each round of boosting.
4 Density Estimation Trees and Density Estimation Forests
Density Estimation Trees (DET) were proposed by Ram & Gray (2011) as an alternative to histograms and kernel density estimation but have received little attention as generative models for sampling or other applications. They model the density function as a constant value over the support of each leaf in a binary tree,
$q(\mathbf{x})=\sum_{j=1}^{J}\frac{\hat{P}(X_{j})}{V(X_{j})}\mathbf{1}_{X_{j}}(\mathbf{x})\;,$ (13)
with $\hat{P}(X)$ being an empirical estimate of probability of the event $\mathbf{x}\in X$ and $V(X)$ denoting the volume of $X$. Note that it is possible to draw an exact sample from this type of model by randomly selecting a leaf, $j\in[1..J]$, given probabilities $\hat{P}(X_{j})$, and then drawing a sample from a uniform distribution over $X_{j}$.
To fit a DET, Ram & Gray (2011) propose optimizing the Integrated Squared Error (ISE) between the data and model distributions which, following a similar approach to Section 3.1, leads the following optimization problem when considering how to split a leaf node:
$\max_{X_{L},X_{R}\in\mathcal{P}\left(X_{P}\right)}D(P(X_{L}),V(X_{L}))+D(P(X_{R}),V(X_{R}))-D(P(X_{P}),V(X_{P}))\,.$ (14)
For the ISE, $D$ should be taken as the function $D_{ISE}(P,V)=\nicefrac{{P^{2}}}{{V}}$ which leads to a similar splitting criterion to Equation 14 but replacing the previous model’s distribution with the volume measure $V$ which can be interpreted as the uniform distribution on $\mathcal{X}$ (up to a multiplicative constant).
Maximum Likelihood
Often generative models are trained to maximize the likelihood of the observed data. This was left for future work in Ram & Gray (2011) but, as we show in Appendix B, can be accomplished by replacing $D_{ISE}$ with $D_{KL}(P,V)=P\,\log\left(\nicefrac{{P}}{{V}}\right)$. This choice of minimization criterion can be seen as analogous to the choice between Gini impurity and Shannon entropy in the computation of the information gain in decision trees.
Bagging and Feature Subsampling
Following the common approach in decision trees, Ram & Gray (2011) suggest the use of pruning for regularization of DET models. Practice has however evolved to prefer bagging as a form of regularization rather than relying on single decision trees. We employ same principle to DETs by fitting many trees on bootstrap samples of the data. We also adopt the common practice from Random Forests of randomly sampling a subset of features to consider when splitting any leaf node in order to encourage independence between the different trees in the ensemble. The ensemble model, which we call Density Estimation Forests (DEF) in the sequel, is thus an additive mixture of DETs with uniform weights, therefore still allowing for normalized density computation and exact sampling.
@@ 5 Related Work
Most prior work on generative boosting focuses on unstructured data and the use of parametric weak learners and is split between two approaches: (i) Additive methods that model the density function as an additive mixture of weak learners such as Rosset & Segal (2002); Tolstikhin et al. (2017). (ii) Those that take a multiplicative approach modeling the density function as an unnormalized product of weak learners. The latter is equivalent to the energy based approach that writes the energy function (log density) as an additive sum of weak learners. Welling et al. (2002) in particular also approach boosting from the point of view of functional optimization of the likelihood or the logistic loss of an energy-based model. However, they rely on a first order local approximation of the objective since they focus on parametric weak learners such as restricted Boltzmann machines for which a second order step would be impractical.
Greedy Multiplicative Boosting
@@
Another more direct multiplicative boosting framework was first proposed by Tu (2007). At each boosting round a discriminative classifier is trained to distinguish between empirical data and data generated by the current model by estimating the likelihood ratio $\nicefrac{{p(\mathbf{x})}}{{q_{t}(\mathbf{x})}}$. This estimated ratio is used as a direct multiplicative factor to update the current model $q_{t}$ (after being raised to an appropriate step size). In ideal conditions this greedy procedure would converge in a single iteration if a step size of 1 would be used. While Tu (2007) does not prescribe a particular choice of classifier, Grover & Ermon (2018) proposes a similar concept where the ratio is estimated based on an adversarial bound for an $f$-divergence and Cranko & Nock (2019) provides additional analysis on this method. We note that the main difference between this greedy approach and NRGBoost is that the latter attempts to update the current density proportionally to an exponential of the likelihood ratio, $\exp\left(\alpha_{t}\cdot\nicefrac{{p(x)}}{{q_{t}(x)}}\right)$, instead of $\left(\nicefrac{{p(\mathbf{x})}}{{q_{t}(\mathbf{x})}}\right)^{\alpha_{t}}$ directly. In Appendix C we explore the differences between NRGBoost and this approach when it is adapted to use trees as weak learners.
Tree-Based Density Modelling
@@
Other authors have proposed tree-based density models similar to DET (Nock & Guillame-Bert, 2022) or additive mixtures of tree-based models (Correia et al., 2020; Wen & Hang, 2022; Watson et al., 2023) but perhaps surprisingly, the natural idea of creating an ensemble of DET models through bagging has not been explored before as far as we are aware. Distinguishing features of some of these alternative approaches are: (i) Not relying on a density estimation goal to drive the partitioning of the input space. Correia et al. (2020) leverages a standard discriminative Random Forest, therefore giving special treatment to a particular input variable whose conditional estimation drives the choice of partitions and Wen & Hang (2022) proposes using a mid-point random tree partitioning. (ii) Leveraging more complex models for the data in the leaf of a tree instead of a uniform density (Correia et al., 2020; Watson et al., 2023). This can allow for the use of trees that are more representative with a smaller number of leaves. (iii) Relying on an adversarial framework where the generator and discriminator are both a tree (Nock & Guillame-Bert, 2022) or an ensemble of trees (Watson et al., 2023). While this also leads to an iterative approach, unlike with boosting or bagging, the new model trained at each round doesn’t add to the previous one but replaces it instead.
Other Recent Tree-Based approaches
@@
Nock & Guillame-Bert (2023) proposes a different ensemble approach where each tree does not have its own leaf values that get added or multiplied to produce the final density, but instead serve to collectively define the partitioning of the input space. To train such models the authors propose a framework where, rather than adding a new tree to the ensemble at every iteration, the model is initialized with a fixed number of tree root nodes and each iteration adds a split to an existing leaf node. Finally Jolicoeur-Martineau et al. (2024) propose a diffusion model where a tree-based model (e.g., GBDT) is used to regress the score function. Being a diffusion model, however, means that computing densities is untractable.
@@ 6 Experiments
We evaluate the our proposed methods on 5 tabular datasets from the UCI Machine Learning Repository (Dheeru & Karra Taniskidou, 2017): Abalone (AB), Physicochemical Properties of Protein Tertiary Structure (PR), Adult (AD), MiniBooNE (MBNE) and Covertype (CT) as well as the California Housing (CH) dataset available through scikit-learn (Pedregosa et al., 2011). We also include a downsampled version of MNIST (by 2x along each dimension) which allows us to visually assess the quality of individual samples, something that is generally difficult with structured tabular data. More details about these datasets are given in Appendix D.1.
We split our experiments into two sections, the first to evaluate the quality of density models directly on a single variable inference task and the second to investigate the performance of our proposed models when used for sampling, comparing them against more specialized models.
6.1 Single Variable Inference
We test the ability of a generative model, trained to learn the density over all input variables, $q(\mathbf{x})$, to infer the value of a single one (i.e., we evaluate the quality of its estimate of $q(x_{i}|\mathbf{x}_{-i})$). For this purpose we pick $x_{i}=y$ as the original target of the dataset, noting that the models that we train do not treat this variable in any special way, except for the selection of the best model in validation. As such, we would expect that the model’s performance in inference over this particular variable is indicative of its strength on any other single variable inference task and also indicative of the quality of the full $q(\mathbf{x})$ from which the conditional probability estimate is derived.
We use XGBoost (Chen & Guestrin, 2016) as a baseline for what should be achievable by a strong discriminative model, noting that it is trained to maximize the discriminative likelihood, $\mathbb{E}_{\mathbf{x}\sim p}\log q(x_{i}|\mathbf{x}_{-i})$, directly, not wasting model capacity in learning other aspects of the full data distribution. We also compare to two other tree-based generative baselines: RFDE (Wen & Hang, 2022) and ARF (FORDE) (Watson et al., 2023). The former allows us to gauge the impact of the guided partitioning used in DEF models over a random partitioning of the input space.
We use random search to tune the hyperparameters of the XGBoost model and a grid search to tune the most important hyperparameters of each generative density model. We employ 5-fold cross-validation, repeating the hyperparameter tuning on each fold. For the full details of the experimental protocol please refer to Appendix D.
@@ 6.1 Single Variable Inference
  • Due to the discrete nature of the MNIST dataset and the fact that the ARF algorithm tries to fit continuous distributions at the leaves, we could not obtain a reasonable density model for this dataset.
We find that NRGBoost outperforms the remaining generative models (see Table 1), even achieving comparable performance to XGBoost on the smaller datasets and with a small gap on the three larger ones (MBNE, MNIST, CT). We note also that for the regression datasets, the generative models provide an estimate of the full conditional distribution over the target variable rather than a point estimate like XGBoost. While there are other variants of boosting that can do the same (Duan et al., 2020), they rely on a parametric assumption about $p(y|\mathbf{x})$ that needs to hold for any $\mathbf{x}$.
@@ 6.2 Sampling
We evaluate two different aspects of the quality of generated samples: their utility for training a machine learning model and how distinguishable they are from real data.
Besides ARF, we compare to TVAE (Xu et al., 2019) and TabDDPM (Kotelnikov et al., 2023), two neural-network based generative models, as well as Forest-Flow (Jolicoeur-Martineau et al., 2024), a tree-based diffusion model. Note that these three methods are not capable of density estimation.
@@
Machine learning (ML) efficiency has been a popular way to measure the quality of generative models for sampling (Xu et al., 2019; Borisov et al., 2022b; Kotelnikov et al., 2023). It relies on using samples from a generative model to train a discriminative model which is then evaluated on real data, and is thus similar to the single variable inference performance that we use to compare density models in Section 6.1. In fact, if the density model’s support covers that of the full data, one would expect the discriminative model to recover the generator’s $q(y|\mathbf{x})$, and therefore its performance, in the limit where infinite generated data is used to train it.
We use an XGBoost model as the discriminative model and train it on the same number of training samples as the original data. For the density models, we generate samples from the best model found in the previous section and for TVAE and TabDDM we select their hyperparameters by evaluating the ML Efficiency in the real validation set (full details of the hyperparameter tuning are provided in Appendix D.2). Note that this leaves these models at a potential advantage since the hyperparameter selection is based on the metric that is being evaluated.
We repeat all experiments 5 times, with different generated datatsets from each model and report the performance of the discriminative model in Table 2. We find that NRGBoost and TabDDPM alternate as the best-performing model depending on the dataset (with two inconclusive cases), and NRGBoost is never ranked lower than second on any dataset. We note also that, despite our best efforts with additional manual tuning, we could not achieve a reasonable model with TabDDM on MNIST.
@@
Similar to Borisov et al. (2022b) we test the capacity of a discriminative model to distinguish between real and generated data. We use the original validation set as the real part of the training data for the discriminator in order to avoid benefiting generative methods that overfit their original training set. A new validation set is carved out of the original test set (20%) and used to tune the hyperparameters of an XGBoost model which we use as our discriminator. We report the AUC of this model on the remainder of the real test data in Table 3 which shows that NRGBoost outperforms other methods except on the PR (with an inconclusive result) and MBNE datasets.
Table 3: AUC of an XGBoost model trained to distinguish real from generated data (lower means better). Reported values are the averages and standard errors over 5 synthetic datasets generated by the same model. The best generative method for each dataset is highlighted in bold and methods that are not significantly worse (as determined by a $t$-test at a 95% confidence level) are underlined.
@@
Statistical Distance Measure
As an additional measure of the statistical dissimilarity between data and model distributions we evaluate a Wasserstein distance using a similar setup to Jolicoeur-Martineau et al. (2024). Namely, we min-max scale the numerical variables, and one-hot encode the categorical variables and scale them by 1/2. We use a $L_{1}$ distance in the formulation of the optimal transport problem. Since finding the optimal solution has a worst case cubic scaling with the sample size, we sub-sample a maximum of 5000 samples from the original train or test sets and use an equal number of synthetic samples. We repeat the evaluation for each method and dataset 5 times using different synthetic data and also different subsampling seeds for the real data (where applicable).
AB CH PR AD MBNE MNIST CT
AB CH PR AD MBNE MNIST CT
Our results for the distance between empirical training and test distribution and synthetic samples from each model are reported in Tables 5 and 4 respectively. We found these results to be somewhat sensitive to the choice of normalization of the numericals and categoricals, as well as to the choice of distance in this normalized space. Despite NRGBoost not being as dominant as in the discriminator measure, it still fares well, obtaining the best average rank over all datasets and experiments. Qualitatively speaking, we find that the samples generated by NRGBoost also look visually similar to the real data in both the MNIST and California datasets (see Figures 1 and 3), being the only generative model that can claim that for the former.

### v2

@@ 3.3 Regularization
The simplest way to regularize a boosting model is to stop training when overfitting is detected by monitoring a suitable performance metric on a validation set. For NRGBoost this could be the increase in log-likelihood at each boosting round (see Appendix A.1). However, estimating this quantity would require drawing additional validation samples from the model. A viable alternative validation strategy which needs no additional samples is to simply monitor a discriminative performance metric. This amounts to monitoring the quality of $q_{f}(x_{i}|\mathbf{x}_{-i})$ instead of the full $q_{f}(\mathbf{x})$.
Besides early stopping, the binary trees themselves can be regularized by limiting the depth or total number of leaves in each tree. Additionally, we can rely on other strategies, such as disregarding splits that would result in a leaf with too little training data, $P(X)$, model data, $Q_{f}(X)$, volume, $V(X)$, or too high of a ratio between training and model data, $\nicefrac{{P(X)}}{{Q_{f}(X)}}$. We found the latter to be the most effective of these, not only yielding better generalization performance than other approaches, but also having the added benefit of allowing us to lower bound the acceptance probability of our rejection sampling scheme. Furthermore, as we show in Appendix A.2, limiting this ratio guarantees that a small enough step size produces an increase in training likelihood at each boosting round.
@@ 4 Related Work
Most prior work on generative boosting focuses on unstructured data and the use of parametric weak learners and is split between two approaches: (i) Additive methods that model the density function as an additive mixture of weak learners such as Rosset & Segal (2002); Tolstikhin et al. (2017). (ii) Those that take a multiplicative approach, modeling the density function as an unnormalized product of weak learners. The latter is equivalent to the energy-based approach that writes the energy function (log-density) as an additive sum of weak learners. Welling et al. (2002), in particular, also approaches boosting from the point of view of functional optimization of the log-likelihood or the logistic loss of an energy-based model. However, it relies on a first order local approximation of the objective given that it focuses on parametric weak learners, such as restricted Boltzmann machines, for which a second-order step would be impractical.
@@
Another more direct multiplicative boosting framework was first proposed by Tu (2007). At each boosting round, a discriminative classifier is trained to distinguish between empirical data and data generated by the current model, by estimating the likelihood ratio $\nicefrac{{p(\mathbf{x})}}{{q_{t}(\mathbf{x})}}$. This estimated ratio is used as a direct multiplicative factor to update a current model, $q_{t}$. In ideal conditions, this greedy procedure would converge in a single iteration (when using a step size of 1). While Tu (2007) does not prescribe a particular choice of classifier, Grover & Ermon (2018) proposes a similar concept where the ratio is estimated based on a variational bound for an $f$-divergence and Cranko & Nock (2019) provides additional analysis on this method. We note that the main difference between this greedy approach and NRGBoost is that the latter attempts to update the current density proportionally to an exponential of the likelihood ratio, $\exp\left(\alpha_{t}\cdot\nicefrac{{p(x)}}{{q_{t}(x)}}\right)$, instead of $\left(\nicefrac{{p(\mathbf{x})}}{{q_{t}(\mathbf{x})}}\right)^{\alpha_{t}}$ directly. In Appendix C we explore the differences between NRGBoost and this approach, when it is adapted to use trees as weak learners.
Tabular Energy-Based Models
Ma et al. (2024) proposes reinterpreting the logits of a TabPFN classifier (Hollmann et al., 2022) as an energy function over the input space. Similarly, Margeloiu et al. (2024) relies on auxiliary binary classification tasks, reinterpreting the logits of the classifiers as class-conditional energy functions. These approaches, while not as principled as maximum likelihood, avoid the need for MCMC sampling and were shown to work well empirically. However, their reliance on TabPFN can limit their applicability to larger datasets. Furthermore, both approaches rely on having a categorical target variable, which further limits their applicability and could introduce an inherent bias given that this variable is not treated in the same way as the others.
Tree-Based Density Models
Density Estimation Trees (DET) were proposed by Ram & Gray (2011) as an alternative to histograms and kernel density estimation. They model the density function as a constant value over the support of each leaf in a binary tree, $q(\mathbf{x})=\sum_{j=1}^{J}\frac{\hat{P}(X_{j})}{V(X_{j})}\mathbf{1}_{X_{j}}(\mathbf{x})$, with $V(X)$ denoting the volume of $X$. The partitioning of the input space is determined directly by greedy optimization of a generative objective (ISE). Despite allowing for efficient exact sampling, DETs have received little attention as generative models. In Appendix B, we explore extensions of this method to maximum likelihood estimation. In our experiments, we also test the natural idea to ensemble DET models through bagging, leading to a generative algorithm analogous to Random Forests.
@@
Other authors have proposed similar tree-based density models (Nock & Guillame-Bert, 2022) or additive mixtures of tree-based models (Correia et al., 2020; Wen & Hang, 2022; Watson et al., 2023). Distinguishing features of some of these alternative approaches are: (i) Not relying on a density estimation goal to drive the partitioning of the input space. Correia et al. (2020) leverages a standard discriminative Random Forest, therefore giving special treatment to a particular input variable whose conditional estimation drives the choice of partitions and Wen & Hang (2022) proposes using a mid-point random tree partitioning. (ii) Leveraging more complex models for the data in the leaf of a tree instead of a uniform density (Correia et al., 2020; Watson et al., 2023). This approach can allow for the use of trees that are more representative with a smaller number of leaves. (iii) Nock & Guillame-Bert (2022) and Watson et al. (2023) both propose generative adversarial frameworks where the generator and discriminator are both a tree or an ensemble of trees respectively. While this also leads to an iterative approach, unlike with boosting or bagging the new model trained at each round doesn’t add to the previous one but replaces it instead.
@@
Nock & Guillame-Bert (2023) proposes a different ensemble approach where trees do not have their own leaf values that are added or multiplied to produce a final density, but instead collectively define the partitioning of the input space. To train such models the authors propose a framework where, rather than adding a new tree to the ensemble at every iteration, the model is initialized with a fixed number of tree root nodes and each iteration adds a split to an existing leaf node. Jolicoeur-Martineau et al. (2024) proposes diffusion and conditional flow matching approaches where a tree-based model (e.g., GBDT) is used to learn the score function or vector field. However, these approaches do not support efficient density estimation. McCarter (2024) leverages the autoregressive framework of TabMT (Gulati & Roysdon, 2024), replacing transformers with XGBoost based predictors. While this approach is capable of density estimation, it requires training one predictor (consisting of one or more XGBoost models) per input variable. Furthermore, these models are trained on datasets created from multiple copies of the original dataset with different missing data patterns, which could limit their applicability to larger datasets.
@@ 5 Experiments
We evaluate NRGBoost on five tabular datasets from the UCI Machine Learning Repository (Dheeru & Karra Taniskidou, 2017): Abalone (AB), Physicochemical Properties of Protein Tertiary Structure (PR), Adult (AD), MiniBooNE (MBNE) and Covertype (CT) as well as the California Housing (CH) dataset available through scikit-learn (Pedregosa et al., 2011). We also include a downsampled version of MNIST (by 2x along each dimension), which allows us to visually assess the quality of individual samples, something that is generally difficult with structured tabular data. More details about these datasets are given in Appendix D.1.
We split our experiments into two sections, the first to evaluate the quality of density models directly on a single-variable inference task and the second to compare the performance of our proposed model on sampling use cases against more specialized models.
5.1 Single-Variable Inference
We test the ability of a generative model, trained to learn the density over all input variables, $q(\mathbf{x})$, to infer the value of a single one given the others. Specifically, we evaluate the quality of its estimate of $q(x_{i}|\mathbf{x}_{-i})$. For this purpose we pick $x_{i}=y$ as the original target of the dataset, noting that the models that we train do not treat this variable in any special way, except when selecting the best model in validation. As such, we would expect that a model’s performance in inference over this particular variable is indicative of its strength on any other single-variable inference task and also indicative of the quality of the full $q(\mathbf{x})$, from which the conditional probability estimate is derived.
We use XGBoost (Chen & Guestrin, 2016) as a baseline for what should be achievable by a strong discriminative model. Note that this model is trained to maximize the discriminative likelihood, $\mathbb{E}_{\mathbf{x}\sim p}\log q(x_{i}|\mathbf{x}_{-i})$, not wasting model capacity in learning other aspects of the full data distribution. Because generative models provide an estimate of the full conditional distribution for the regression datasets, rather than a single point estimate like XGBoost, we also include NGBoost (Duan et al., 2020) as another discriminative baseline that can estimate the full $q(x_{i}|\mathbf{x}_{-i})$ on these datasets. Note, however, that NGBoost relies on a parametric assumption about this conditional distribution, that needs to hold for any $\mathbf{x}_{-i}$. We use the default assumption of a Normal distribution for all datasets.
We compare NRGBoost against a bagging ensemble of DET models (Ram & Gray, 2011), which we call Density Estimation Forests (DEF), trained for minimizing either ISE or KL divergence (see Appendix B). We also include two other tree-based generative baselines: RFDE (Wen & Hang, 2022) and ARF (Watson et al., 2023). The former allows us to gauge the impact of the guided partitioning used in DEF models over a random partitioning of the input space.
We use random search to tune the hyperparameters of XGBoost and NGBoost and a grid search to tune the most important hyperparameters of each generative density model. We employ 5-fold cross-validation, repeating the hyperparameter tuning on each fold. For the full details of the experimental protocol please refer to Appendix D.
@@ 5.1 Single-Variable Inference
  • Due to the discrete nature of the MNIST dataset and the fact that the ARF algorithm tries to fit continuous distributions for numerical variables at the leaves, we could not obtain a reasonable density model for this dataset.
We find that NRGBoost outperforms the remaining generative models (see Table 1), even achieving comparable performance to XGBoost on the smaller datasets and with a small gap on the three larger ones (MBNE, MNIST and CT). Furthermore, in Appendix E.3 we show that, for the same tasks but in the presence of a missing covariate, NRGBoost can outperform XGBoost relying on popular imputation strategies due to its principled handling of missing data through marginalization.
@@ 5.2 Sampling
We evaluate two different aspects of the quality of generated samples: their utility for training a machine learning model and how distinguishable they are from real data. In addition to the previous ARF and DEF (KL) baselines, we compare against TVAE (Xu et al., 2019) and TabDDPM (Kotelnikov et al., 2023), two deep-learning-based generative models, as well as Forest-Flow (Jolicoeur-Martineau et al., 2024), a tree-based conditional flow matching model.
@@
Machine learning (ML) efficiency has been a popular way to measure the quality of generative models for sampling (Xu et al., 2019; Kotelnikov et al., 2023; Borisov et al., 2022b). It relies on using samples from the model to train a discriminative model which is then evaluated on real data, and is thus similar to the single-variable inference performance that we use to compare density models in Section 5.1. In fact, if the discriminative model is flexible enough, one would expect it to recover the generator’s $q(y|\mathbf{x})$, and therefore its performance, in the limit where infinite generated data is used to train it.
We use XGBoost as the discriminative model and train it on the same number of synthetic samples as in the original training data. For the density models, we generate the samples from the best model found in the previous section. For TVAE and TabDDPM, we select their hyperparameters by evaluating the ML efficiency in the real validation set (full details of the hyperparameter tuning are provided in Appendix D.3). Note that this leaves these models at a potential advantage since the hyperparameter selection is based on the metric that is being evaluated.
We repeat all experiments 5 times, with different generated datatsets from each model and report the performance of the discriminative model in Table 2. We find that NRGBoost and TabDDPM alternate as the best-performing model depending on the dataset (with two inconclusive cases), and NRGBoost is never ranked lower than second on any dataset. We also note that, despite our best efforts with additional manual tuning, we could not obtain a reasonable model with TabDDPM on MNIST.
@@
Similar to Borisov et al. (2022b) we test the capacity of a discriminative model to distinguish between real and generated data. We use the original validation set as the real part of the training data for the discriminator in order to avoid benefiting generative methods that overfit their original training set. A new validation set is carved out of the original test set (20%) and used to tune the hyperparameters of an XGBoost model which we use as our discriminator. We report the AUC of this model on the remainder of the test data in Table 3, which shows that NRGBoost outperforms other methods except on the PR (where the result is inconclusive) and MBNE datasets.
In Appendix E.1 we report a Wasserstein distance estimate computed using a similar setup to Jolicoeur-Martineau et al. (2024) as an alternative measure of the statistical distance between real and model distributions. This evaluation also shows NRGBoost as ranking the best on average across datasets. Qualitatively speaking, its samples also look visually similar to the real data in both the California and MNIST datasets (see Figures 3 and 4).
Figure 4: Downsampled MNIST samples generated by the best generative models on this dataset. Despite being a simple dataset that would pose no challenges to image models, it is hard for tabular generative models due to the high dimensionality and complex structure of correlations between features. We find NRGBoost to be the only tabular model that is able to generate passable samples.
Table 3: AUC of an XGBoost model trained to distinguish real from generated data (lower is better). Reported values are the averages and standard errors over 5 synthetic datasets generated by the same model. The best generative method for each dataset is highlighted in bold and methods that are not significantly worse (as determined by a $t$-test at a 95% confidence level) are underlined.